OpenStudy (anonymous):

10 + -6 ÷ 2(9 – -6^2)

OpenStudy (anonymous):

9.9333333333 which is (149/15)

OpenStudy (dumbcow):

if you divide first 10 + -3*(45) =-125

OpenStudy (anonymous):

i dont understand how you got that answer

OpenStudy (dumbcow):

10 + -6 / 2*(45) 10 + -3*45 10 - 135 -125

OpenStudy (anonymous):

i forgot the negative on the ten is that going to make a difference

jhonyy9 (jhonyy9):

hence you think -10 + -6 +2(9- -6*2) = -10 -6 +2(9-36)=-16+2(-27)=-16-54=-70